A gallon of oil is poured to a cubical container 7 inches on an edge. About how high in inches does the oil rise in the container? (I gallon = 231 cubic inches)
Answer by mananth(16949)

volume of cube = l*l*l
l=7
let height of oil be h
1 gallon = 231 in^3
231= 7*7*h
231/49 = h
4.71 inches is the height of oil